# Inference on many jumps in nonparametric panel regression models

### v2

@@ 4.1 Data Generating Processes
$Y_{jt}=\cos(X_{jt})+\sin(U_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq 0\}}+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt},$ (4.1)
@@
$\sigma_{j}(x,u)=1+\left(\frac{3}{8}-\frac{\left|x\right|}{4}\right)\left(\frac{3}{2}\right)^{2u}.$ (4.2)
  1. DGP1:
    i.i.d. setting with heteroskedasticity. $X_{jt}\sim i.i.d.\ U[-1,1]$, $\epsilon_{jt}\sim i.i.d.\ N(0,1)$.
  2. DGP2:
@@
    Factor structure and homoskedasticity. Both the running variable and error term have a factor structure, $\epsilon_{jt}=\lambda_{\epsilon,j}f_{\epsilon,t}+u_{\epsilon,jt}$ and $X_{jt}=1/4\left(\lambda_{x,j}f_{x,t}+u_{x,jt}\right)$, where $f_{\epsilon,t}$, $u_{\epsilon,jt}$, $f_{x,t}$ and $u_{x,jt}$ are each generated from an MA($\infty$) process with algebraic decay parameter, $\beta=1.5$. The factor loadings, $\lambda_{\epsilon,j}$ and $\lambda_{x,j}$, are i.i.d. draws from a standard normal distribution.
  3. DGP3:
@@
    Factor structure and heteroskedasticity. The setting is the same as in DGP2 except that the error term is heteroskedastic with the conditional standard deviation described in (4.2).
@@
All DGPs, except for the first one, exhibit strong cross-sectional and weak temporal correlation in the covariate variable and the error term. We are interested in testing the existence of the jump effect, namely, $H_{0}^{\left(1\right)}:\gamma_{j}=0$ for all $j\in[N]$. For each DGP, we study both the size and local power performance of our testing procedure. When studying the size performance, we set all $\gamma_{j}=0$. When examining the local power, we set a fraction of the threshold effects to be $\gamma_{j}=T^{-2/5}(\mathrm{log}N)^{1/2}B_{j}$, where $B_{j}\sim i.i.d.\ U[2,10]$, and the remaining $\gamma_{j}$’s to be zero. We consider two different choices for the cross-sectional dimension: $N=10,$ 100. When $N=10$, the above fraction is set to $20\%$, whereas in the high-dimensional setting with $N=100,$ the fraction is decreased to $10\%$.
4.2 Simulation Results
Throughout the study, we use a local linear estimator with a uniform kernel and select a mean square error-optimal bandwidth following the procedure described in Calonico et al., (2014). We consider three significance levels, namely, $\alpha=0.1,$ $0.05,$ and $0.01$. The results are based on $1000$ Monte Carlo iterations.
The results of simultaneous jump effects testing for DGP1 through DGP3 are displayed in the Table 1. First, the empirical size closely matches the nominal size across all $N$ values and for large $T$ in all DGPs. For DGP3, we observe slightly oversize when $N=10$ which may be caused by the presence of heteroskedasticity in comparison with results with DGP2. Overall, these simulation results confirm our theoretical findings in Theorem 3.1. In particular, it shows that the test procedure, which is based on critical values neglecting the dependence in the covariance structure, has fairly accurate size control even in the case of strong cross-sectional dependence in the innovations. Second, the local power properties of the test are reasonably well for all DGPs under investigation, and the power increases as either $N$ or $T$ increases. This result confirms that our testing procedure is well-suited to detect sparse alternatives.
In addition, we present comprehensive simulation results in Section D in the supplementary material. First, we achieve comparable simulation performance in terms of size and power results for the test of homogeneity of the threshold effects in Table A2. The results confirm our Gaussian approximation result in Theorem 3.2. For robustness, we also consider three additional settings with even stronger cross-sectional dependence, see DGP4–6 in Section D of the supplementary materials. The results in Table A1 show that size and power are indeed robust to strong cross-sectional dependence. Furthermore, Table A3 demonstrates that detecting the existence of jump effects simultaneously across cross-sectional dimensions in the case of unknown thresholds has size and power properties as good as in the known threshold case, confirming our theoretical results from Theorem 3.5. Finally, we show the good finite sample performance for the estimation accuracy of the threshold location estimator in Table A4, providing Monte Carlo evidence supporting the consistency result of the threshold in Theorem 3.6.
To demonstrate the practical relevance of our testing procedure, we present two real data applications in Section E of the supplementary material. In Subsection E.1, we detect significant threshold effects in the news impact curve (Engle and Ng, , 1993), i.e., in a model where past returns influence present volatility of stocks. It is an illustration of our testing procedure under unknown thresholds. As a second application, we revisit the study of the incumbency effect in US House elections in Subsection E.2, and find significant jump effects as well as heterogeneity across different states. Since the threshold is naturally given in this problem, it functions as an illustration of our testing procedures for the existence of jump effects and the test for homogeneity in the known case.

### v3

@@ E.1 Data Generating Processes
$Y_{jt}=\cos(X_{jt})+\sin(U_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq 0\}}+\sigma_{j}(X_{jt},U_{jt})\epsilon_{jt},$ (SA-1)
@@
$\sigma_{j}(x,u)=1+\left(\frac{3}{8}-\frac{\left|x\right|}{4}\right)\left(\frac{3}{2}\right)^{2u}.$ (SA-2)
  1. DGP1: i.i.d. setting with heteroskedasticity. $X_{jt}\sim i.i.d.\ U[-1,1]$, $\epsilon_{jt}\sim i.i.d.\ N(0,1)$.
@@
  2. DGP2: Factor structure and homoskedasticity. Both the running variable and error term have a factor structure, $\epsilon_{jt}=\lambda_{\epsilon,j}f_{\epsilon,t}+u_{\epsilon,jt}$ and $X_{jt}=1/4\left(\lambda_{x,j}f_{x,t}+u_{x,jt}\right)$, where $f_{\epsilon,t}$, $u_{\epsilon,jt}$, $f_{x,t}$ and $u_{x,jt}$ are each generated from an MA($\infty$) process with algebraic decay parameter, $\beta=1.5$. The factor loadings, $\lambda_{\epsilon,j}$ and $\lambda_{x,j}$, are i.i.d. draws from a standard normal distribution.
@@
  3. DGP3: Factor structure and heteroskedasticity. The setting is the same as in DGP2 except that the error term is heteroskedastic with the conditional standard deviation described in (SA-2).
@@
Additionally, we examine three additional cases, referred to as DGP4–6 in the Supplementary Material E, which account for varying degrees of cross-sectional dependence and a possible time effect. All DGPs, except for the first one, exhibit strong cross-sectional and weak temporal correlation in the covariate variable and the error term. We are interested in testing the existence of the jump effect, namely, $H_{0}^{\left(1\right)}:\gamma_{j}=0$ for all $j\in[N]$. For each DGP, we study both the size and local power performance of our testing procedure. When studying the size performance, we set all $\gamma_{j}=0$. When examining the local power, according to Theorem A.3 in the Supplement and to the selection of the bandwidth, we set a fraction of the threshold effects to be $\gamma_{j}=T^{-2/5}(\mathrm{log}N)^{1/2}B_{j}$, where $B_{j}\sim i.i.d.\ U[2,10]$, and the remaining $\gamma_{j}$’s to be zero. We consider two different choices for the cross-sectional dimension: $N=10,$ 100. When $N=10$, the above fraction is set to $20\%$, whereas in the high-dimensional setting with $N=100,$ the fraction is decreased to $10\%$.
E.2 Additional Simulation Results